# Universal Discrete-Domain Speech Enhancement

### v1

@@ IV-A Dataset Construction and Task Definition
We constructed clean speech dataset from the VoiceBank corpus [57], with 23,075 utterances from 56 speakers for training set and 824 utterances from 2 unseen speakers for test set. The speech sampling rate is 44.1 kHz. We constructed task-specific degraded datasets based on a clean speech dataset, incorporating three conventional distortions (i.e., noise, reverberation, and band-limiting), three unconventional distortions (i.e., clipping, phase distortion, and compression distortion), and three mixed distortions. Each task is defined as follows.
1) Denoising (DN): This task aims to enhance the degraded speech caused by additive noise. We constructed the noisy dataset by adding noise from the DEMAND dataset [58] to the clean speech. For the training set, we used 5 types of noise, with the signal-to-noise ratio (SNR) ranging from 0 to 15 dB, in 5 dB intervals. For the test set, we used 5 types of unseen noise, with the SNR ranging from 2.5 to 17.5 dB in 5 dB intervals.
2) Dereverberation (DR): This task aims to enhance the degraded speech caused by reverberation. We adopted the room impulse response (RIR) dataset from the DNS Challenge [59], which included 248 real RIRs and approximately 60,000 simulated RIRs. We constructed the reverberant dataset by randomly selecting RIRs and convolving RIRs with clean speech. The RIRs of the test set were unseen in the training set.
@@
5) Phase Distortion Restoration (PDR): This task aims to enhance the degraded speech caused by inaccurate phase. We constructed the phase distortion dataset by replacing phase spectra of clean speech with the ones predicted by an NSPP [^1] https://github.com/YangAi520/NSPP. model [29].
6) Compression Distortion Restoration (CDR): This task aims to enhance the degraded speech caused by compression from a speech codec. We constructed the compression distortion dataset by compressing clean speech using an APCodec [^2] https://github.com/YangAi520/APCodec. [46] with a single VQ.
7) Mixed Distortion Restoration: We combined the six types of distortions mentioned above to create three tasks, represented as DN+DR+BWE, which stands for denoising + dereverberation + bandwidth extension (band-limited speech sampling rate is 8 kHz), DN+DR+DC, which stands for denoising + dereverberation + declipping, and DN+PDR+CDR, which stands for denoising + phase distortion restoration + compression distortion restoration.
@@
DEMUCS 3.49 4.03 3.19 3.11 3.79
CMGAN 3.49 4.03 3.20 4.57 4.04
MP-SENet 3.51 4.04 3.22 4.53 4.08
UDSE 3.50 4.03 3.21 4.49 4.03
@@
DEMUCS 1.86 3.51 1.51 0.78 1.34
CMGAN 2.05 3.39 1.73 1.04 1.34
MP-SENet 3.11 3.94 2.79 1.54 1.93
UDSE 3.43 4.01 3.13 4.25 3.47
@@ IV-B Model Details
The UDSE [^3] Speech samples are available at: https://anonymity225.github.io/UDSE/. utilized a DAC [^4] https://github.com/descriptinc/descript-audio-codec. [45] as $\phi_{Codec}$, which included 9 VQs (i.e., $N=9$), with both the codebook size and the code vector dimension set to 1024 (i.e., $K=M=1024$). The feature processor $\phi_{FP}$ used 8 Conformer blocks (i.e., $B_{G}=8$), while each token predictor $\phi_{TP_{1}},\dots,\phi_{TP_{N}}$ used 4 Conformer blocks (i.e., $B_{T}=4$). The number of channels and attention heads in each block were 512 and 8, respectively (i.e., $C=512$). We trained the UDSE using an AdamW optimizer on a single Nvidia A800 GPU, with $\beta_{1}=0.9,\beta_{2}=0.95,$ and a weight decay of 0.01 for 100 epochs. The initial learning rate was set to 0.0005, with a cosine annealing strategy for decay and a warm-up training scheduler for the first 4k steps.
We compared UDSE with several advanced continuous-domain SE model, including waveform-modeling-based DEMUCS [^5] https://github.com/facebookresearch/denoiser. [17], and spectrum-modeling-based CMGAN [^6] https://github.com/ruizhecao96/CMGAN. [25] and MP-SENet [^7] https://github.com/yxlu-0102/MP-SENet. [26, 27]. For task BWE, we also compared UDSE with the advanced AP-BWE [^8] https://github.com/yxlu-0102/AP-BWE. [60], specifically designed for speech bandwidth extension. We reproduced these models using the official codes they provided.
TABLE II: Average preference scores (%) of ABX subjective tests on speech quality between UDSE and a baseline (i.e., model B) on different SE tasks, where N/P stands for “no preference” and $p$ denotes the $p$-value of a $t$-test between two models. For tasks DN, DR, DC, PDR, DN+DR+BWE, DN+DR+DC and DN+PDR+CDR, “model B” is MP-SENet, while for tasks BWE and CDR, “model B” is AP-BWE and CMGAN, respectively.
@@ IV-C Evaluation Metrics
For objective evaluation, since UDSE predicts the distribution of discrete tokens corresponding to clean speech (i.e., it formulates SE as a classification task rather than directly generating the clean speech as in regression-based models), traditional intrusive indicators such as perceptual evaluation of speech quality (PESQ) are not well-suited for evaluating our model, as suggested in [34]. Therefore, we used three non-intrusive metrics commonly adopted in SE field, including NISQA [61], DNSMOS [62] and UTMOS [63], to assess overall speech quality. NISQA is a full-band metric that can evaluate the overall quality of 44.1 kHz speech, while DNSMOS and UTMOS both design for 16 kHz, focus more on the low-frequency quality. All three metrics use the same scoring scale as traditional mean opinion score (MOS), ranging from 1 to 5. Both NISQA and UTMOS directly reflect the perceived acceptability and naturalness of speech. DNSMOS, on the other hand, follows the ITU-T P.835 subjective test framework and provides multi-dimensional scores: SIG represents speech quality and primarily reflects overall signal distortion; BAK measures the perceptual intensity of background noise, with higher scores indicating better noise suppression; and OVRL provides an overall rating, indicating the general naturalness of the speech.
@@ IV-D1 Experimental Results for Conventional Distortions
For the classic dereverberation task DR, our proposed UDSE significantly outperformed the advanced regression-based continuous-domain baseline models in most objective metrics, as shown in Table I. Specifically, UDSE significantly outperformed all baselines in both NISQA and UTMOS scores, and achieved a comparable DNSMOS score to MP-SENet, effectively suppressing speech distortion caused by reverberation. It was evident that although CMGAN performed well in denoising tasks, it lagged behind in all objective metrics when dealing with dereverberation. In contrast, the proposed UDSE consistently achieved robust performance across both distortion types. The second row of Figure 3 shows the spectrograms of reverberant speech enhanced by MP-SENet and UDSE under low reverberation time conditions. Interestingly, the spectrogram of the speech enhanced by MP-SENet is more similar to that of the clean reference speech. This may be attributed to the fact that regression-based continuous-domain SE models directly model the speech waveform or features. In particular, MP-SENet explicitly predicts the amplitude spectra, which makes the enhanced speech more “similar” to the reference speech, especially when the distortion is not severe (e.g., under high SNR or low reverberation time conditions). In contrast, classification-based discrete-domain models indirectly predict discrete tokens of speech and then decode them into speech waveforms, which may result in enhanced speech that is “not similar” to the reference speech. This is also why intrusive indicators that rely on reference speech are not suitable for evaluating discrete-domain models, as stated in Section IV-C. At this point, subjective evaluation becomes the gold standard for assessing speech quality. As shown in the results in Table II, this lack of “similarity” does not negatively impact perceptual quality — reverberant speech enhanced by UDSE is even significantly better than that enhanced by MP-SENet ($p<0.01$).
@@
For the bandwidth extension task BWE, as suggested in Table I, DEMUCS, CMGAN and MP-SENet, which were specifically designed for denoising and reverberation, all failed. Although the objective results of AP-BWE, which was specifically designed for task BWE, outperformed the three models mentioned above, there is still a significant gap compared to our proposed UDSE. Subjectively, UDSE also significantly outperformed AP-BWE ($p<0.01$) as shown in Table II. According to listener feedback, the speech extended by AP-BWE contained noticeable pronunciation errors, whereas UDSE did not. The last row of Figure 3 intuitively shows the results of bandwidth extension. It is evident that UDSE achieves a richer and more accurate reconstruction of high-frequency speech details compared to AP-BWE. For example, within the time range of 1 to 1.2 seconds and the frequency band of 1 kHz to 6 kHz, the harmonic details of the speech extended by UDSE are more similar to those of the clean speech. This indicates that UDSE has potential in speech bandwidth extension for extremely narrow bands (e.g., from 2 kHz to 44.1 kHz, representing an expansion of over 22 times in bandwidth).
@@ IV-D2 Experimental Results for Unconventional Distortions
For the phase distortion restoration task PDR, UDSE achieved objective results comparable to those of MP-SENet, while outperforming DEMUCS and CMGAN, as shown in Table I. MP-SENet, by introducing explicit phase prediction, was more suitable for phase distortion recovery compared to DEMUCS and CMGAN. As shown in the second row of Figure 4, speech affected by phase distortion exhibited prominent horizontal stripes — perceptually experienced as an irritating buzzing sound. In contrast, the spectrogram of the UDSE-enhanced speech showed no such patterns, indicating that UDSE effectively eliminated interference caused by phase distortion. Since phase restoration effects are difficult to visualize clearly in spectrograms, subjective evaluation is necessary to provide additional evidence. As shown in the subjective evaluation results in Table II, UDSE-enhanced speech was generally preferred by listeners ($p<0.01$).
For the compression distortion restoration task CDR, our proposed UDSE significantly outperformed all baseline models in both objective and subjective aspects ($p<0.01$), as shown in Tables I and II. Interestingly, we found that these baseline models even had a negative effect on the enhancement of compressed speech. As clearly shown in the last row of Figure 4, UDSE effectively restored spectral details, with the enhanced harmonics closely resembling those of the clean speech. Although the speech enhanced by CMGAN showed decent restoration of the low-frequency harmonics, it lost many spectral details in the high-frequency regions and exhibited numerous unnatural horizontal stripes, which significantly degraded the listening experience. These results suggest that continuous-domain SE models are less effective in handling codec-induced compression distortions, likely due to the strong correlation between compressed and clean speech signals. In comparison, the proposed UDSE tackles this challenge from a discrete-domain perspective.
@@ IV-D3 Experimental Results for Mixed Distortions
In the three mixed distortion restoration tasks, our proposed UDSE had a significant advantage over all baseline models in both objective and subjective aspects ($p<0.01$), as shown in Tables I and II. For task DN+DR+BWE, UDSE outperformed MP-SENet by more than 2 points in the NISQA score. In addition, for task DN+DR+DC, UDSE achieved a 2.7-point higher OVRL score in DNSMOS and a 1.5-point higher UTMOS score compared to MP-SENet. Consistently, for task DN+PDR+CDR, UDSE again exceeded MP-SENet by over 2 points in the NISQA metric. Subjective evaluations showed that, in all three mixed-distortion tasks, the number of participants preferring UDSE-enhanced speech was at least double that of those favoring MP-SENet — highlighting UDSE’s superior performance in managing complex distortions.
@@
The spectrograms shown in Figure 5 further provide a clear visual illustration of the effectiveness of the proposed UDSE in enhancing speech with mixed distortions. We can observe that for the DN+DR+BWE task (the first row in Figure 5), UDSE recovered the high-frequency components much more accurately, bringing the result closer to the ground-truth clean speech, while MP-SENet struggled with recovering the high-frequency part. For the DN+DR+DC task (the second row in Figure 5), the speech enhanced by MP-SENet exhibited horizontal stripes in all-frequency bands, resulting in a mechanical listening experience and a loss of the original speaker’s timbral characteristics. For the DN+PDR+CDR task (the last row in Figure 5), the spectrogram of the speech recovered by UDSE retained more details, while MP-SENet produced a smoothed spectrogram with noticeable noise, significantly impacting the listening experience. All the above experimental results demonstrate that the proposed UDSE can effectively mitigate various types of distortion in degraded speech, making it more suitable for real-world applications.

### v2

@@ IV-A Dataset Construction and Task Definition
We constructed clean speech dataset from the VoiceBank corpus [valentini2016investigating], with 23,075 utterances from 56 speakers for training set and 824 utterances from 2 unseen speakers for test set. The speech sampling rate is 44.1 kHz. We constructed task-specific degraded datasets based on a clean speech dataset, incorporating three conventional distortions (i.e., noise, reverberation, and band-limiting), three unconventional distortions (i.e., clipping, phase distortion, and compression distortion), and three mixed distortions. Each task is defined as follows.
1) Denoising (DN): This task aims to enhance the degraded speech caused by additive noise. We constructed the noisy dataset by adding noise from the DEMAND dataset [thiemann2013diverse] to the clean speech. For the training set, we used 5 types of noise, with the signal-to-noise ratio (SNR) ranging from 0 to 15 dB, in 5 dB intervals. For the test set, we used 5 types of unseen noise, with the SNR ranging from 2.5 to 17.5 dB in 5 dB intervals.
2) Dereverberation (DR): This task aims to enhance the degraded speech caused by reverberation. We adopted the room impulse response (RIR) dataset from the DNS Challenge [dubey2024icassp], which included 248 real RIRs and approximately 60,000 simulated RIRs. We constructed the reverberant dataset by randomly selecting RIRs and convolving RIRs with clean speech. The RIRs of the test set were unseen in the training set.
@@
5) Phase Distortion Restoration (PDR): This task aims to enhance the degraded speech caused by inaccurate phase. We constructed the phase distortion dataset by replacing phase spectra of clean speech with the ones predicted by an NSPP [^1] https://github.com/YangAi520/NSPP. model [ai2023neural].
6) Compression Distortion Restoration (CDR): This task aims to enhance the degraded speech caused by compression from a speech codec. We constructed the compression distortion dataset by compressing clean speech using an APCodec [^2] https://github.com/YangAi520/APCodec. [ai2024apcodec] with a single VQ.
7) Mixed Distortion Restoration: We combined the six types of distortions mentioned above to create three tasks, represented as DN+DR+BWE, which stands for denoising + dereverberation + bandwidth extension (band-limited speech sampling rate is 8 kHz), DN+DR+DC, which stands for denoising + dereverberation + declipping, and DN+PDR+CDR, which stands for denoising + phase distortion restoration + compression distortion restoration. We did not downsample the speech to 2 kHz in the DN+DR+BWE task, as in the standalone BWE task, because further adding noise and reverberation to 2 kHz speech would render the speech nearly unintelligible and the SE solution unrealistically ill-posed.
@@
DEMUCS 3.11 3.49 4.03 3.19 3.79
CMGAN 4.57 3.49 4.03 3.20 4.04
MP-SENet 4.53 3.51 4.04 3.22 4.08
UniverSE++ 3.98 3.48 4.05 3.19 3.94
Genhancer 3.81 3.26 3.96 2.96 3.14
UDSE 4.49 3.50 4.03 3.21 4.03
@@
DEMUCS 0.78 1.86 3.51 1.51 1.34
CMGAN 1.04 2.05 3.39 1.73 1.34
MP-SENet 1.54 3.11 3.94 2.79 1.93
UniverSE++ 3.74 2.92 3.97 2.65 2.36
Genhancer 2.00 2.11 3.62 1.92 1.62
UDSE 4.25 3.43 4.01 3.13 3.47
@@ IV-B Model Details
The UDSE [^3] Codes and speech samples are available at: https://fliu215.github.io/UDSE/. utilized a DAC [^4] https://github.com/descriptinc/descript-audio-codec. [kumar2023high] as $\phi_{Codec}$, which included 9 VQs (i.e., $N=9$), with both the codebook size and the code vector dimension set to 1024 (i.e., $K=M=1024$). The feature processor $\phi_{FP}$ used 8 Conformer blocks (i.e., $B_{G}=8$), while each token predictor $\phi_{TP_{1}},\dots,\phi_{TP_{N}}$ used 4 Conformer blocks (i.e., $B_{T}=4$). The number of channels and attention heads in each block were 512 and 8, respectively (i.e., $C=512$). We trained the UDSE using an AdamW optimizer on a single NVIDIA A800 GPU, with $\beta_{1}=0.9,\beta_{2}=0.95,$ and a weight decay of 0.01 for 100 epochs. The initial learning rate was set to 0.0005, with a cosine annealing strategy for decay and a warm-up training scheduler for the first 4k steps.
We compared UDSE with several advanced continuous-domain SE models, including waveform-regression-based DEMUCS [^5] https://github.com/facebookresearch/denoiser. [defossez2020real], spectrum-regression-based CMGAN [^6] https://github.com/ruizhecao96/CMGAN. [abdulatif2024cmgan] and MP-SENet [^7] https://github.com/yxlu-0102/MP-SENet. [lu2023mp, lu2025explicit], and diffusion-based UniverSE++ [^8] https://github.com/line/open-universe. [universe++]. In addition, we have also reproduced the Genhancer [yang2024genhancer] as a discrete-domain baseline. For task BWE, we also compared UDSE with the advanced continuous-domain AP-BWE [^9] https://github.com/yxlu-0102/AP-BWE. [lu2024towards], specifically designed for speech bandwidth extension. Except for Genhancer, we reproduced all the models using the official codes they provided, whereas for Genhancer, we had to reproduce it manually due to the lack of open-source code. All models were conducted on our 44.1 kHz dataset.
TABLE II: Average preference scores (%) of ABX subjective tests on speech quality between UDSE and a baseline (i.e., model B) on different SE tasks, where N/P stands for “no preference” and $p$ denotes the $p$-value of a $t$-test between two models. For tasks DN, DR, DC and PDR, “model B” is MP-SENet, and for tasks CDR, DN+DR+BWE, DN+DR+DC and DN+PDR+CDR, “model B” is UniverSE++, while for task BWE, “model B” is AP-BWE.
@@ IV-C Evaluation Metrics
For objective evaluation, since UDSE predicts the distribution of discrete tokens corresponding to clean speech (i.e., it formulates SE as a classification task rather than directly generating the clean speech as in regression-based models), traditional intrusive indicators such as perceptual evaluation of speech quality (PESQ) are not well-suited for evaluating our model, as suggested in [yao2025gense]. Therefore, we used three non-intrusive metrics commonly adopted in SE field, including NISQA [mittag2021nisqa], DNSMOS [reddy2022dnsmos] and UTMOS [saeki2022utmos], to assess overall speech quality. NISQA is a full-band metric that can evaluate the overall quality of 44.1 kHz speech, while DNSMOS and UTMOS are both designed for 16 kHz, focus more on the low-frequency quality. All three metrics use the same scoring scale as traditional mean opinion score (MOS), ranging from 1 to 5. Both NISQA and UTMOS directly reflect the perceived acceptability and naturalness of speech. DNSMOS, on the other hand, follows the ITU-T P.835 subjective test framework and provides multi-dimensional scores: SIG represents speech quality and primarily reflects overall signal distortion; BAK measures the perceptual intensity of background noise, with higher scores indicating better noise suppression; and OVRL provides an overall rating, indicating the general naturalness of the speech.
@@ IV-D1 Experimental Results for Conventional Distortions
For the classic dereverberation task DR, our proposed UDSE significantly outperformed the advanced continuous-domain baseline models in most objective metrics, as shown in Table I. Specifically, UDSE significantly outperformed all baselines in both NISQA and UTMOS scores, and achieved a comparable DNSMOS score to MP-SENet, effectively suppressing speech distortion caused by reverberation. It was evident that although CMGAN performed well in denoising tasks, it lagged behind in all objective metrics when dealing with dereverberation. In contrast, the proposed UDSE consistently achieved robust performance across both distortion types. The second row of Figure 3 shows the spectrograms of reverberant speech enhanced by MP-SENet and UDSE under low reverberation time conditions. Interestingly, the spectrogram of the speech enhanced by MP-SENet is more similar to that of the clean reference speech. This may be attributed to the fact that regression-based continuous-domain SE models directly model the speech waveform or features. In particular, MP-SENet explicitly predicts the amplitude spectra, which makes the enhanced speech more “similar” to the reference speech, especially when the distortion is not severe (e.g., under high SNR or low reverberation time conditions). In contrast, classification-based discrete-domain models indirectly predict discrete tokens of speech and then decode them into speech waveforms, which may result in enhanced speech that is “not similar” to the reference speech. This is also why intrusive indicators that rely on reference speech are not suitable for evaluating discrete-domain models, as stated in Section IV-C. At this point, subjective evaluation becomes the gold standard for assessing speech quality. As shown in the results in Table II, this lack of “similarity” does not negatively impact perceptual quality — reverberant speech enhanced by UDSE is even significantly better than that enhanced by MP-SENet ($p<0.01$).
@@
For the bandwidth extension task BWE, as suggested in Table I, DEMUCS, CMGAN, MP-SENet and Genhancer, which were specifically designed for denoising and reverberation, all failed. The more general-purpose UniverSE++ also performed poorly. Although the objective results of AP-BWE, which was specifically designed for task BWE, outperformed the five models mentioned above, there is still a significant gap compared to our proposed UDSE. Subjectively, UDSE also significantly outperformed AP-BWE ($p<0.01$) as shown in Table II. According to listener feedback, the speech extended by AP-BWE contained noticeable pronunciation errors, whereas UDSE did not. The last row of Figure 3 intuitively shows the results of bandwidth extension. It is evident that UDSE achieves a richer and more accurate reconstruction of high-frequency speech details compared to AP-BWE. For example, within the time range of 1 to 1.2 seconds and the frequency band of 1 kHz to 6 kHz, the harmonic details of the speech extended by UDSE are more similar to those of the clean speech. This indicates that UDSE has potential in speech bandwidth extension for extremely narrow bands (e.g., from 2 kHz to 44.1 kHz, representing an expansion of over 22 times in bandwidth).
@@ IV-D2 Experimental Results for Unconventional Distortions
For the phase distortion restoration task PDR, UDSE achieved objective results comparable to those of MP-SENet, while outperforming DEMUCS, CMGAN, UniverSE++ and Genhancer, as shown in Table I. MP-SENet, by introducing explicit phase prediction, was more suitable for phase distortion recovery compared to DEMUCS, CMGAN, UniverSE++ and Genhancer. As shown in the second row of Figure 4, speech affected by phase distortion exhibited prominent horizontal stripes — perceptually experienced as an irritating buzzing sound. In contrast, the spectrogram of the UDSE-enhanced speech showed no such patterns, indicating that UDSE effectively eliminated interference caused by phase distortion. Since phase restoration effects are difficult to visualize clearly in spectrograms, subjective evaluation is necessary to provide additional evidence. As shown in the subjective evaluation results in Table II, UDSE-enhanced speech was generally preferred by listeners ($p<0.01$).
For the compression distortion restoration task CDR, our proposed UDSE significantly outperformed all baseline models in both objective and subjective aspects ($p<0.01$), as shown in Tables I and II. Interestingly, we found that these baseline models even had a negative effect on the enhancement of compressed speech. As clearly shown in the last row of Figure 4, UDSE effectively restored spectral details, with the enhanced harmonics closely resembling those of the clean speech. Although the speech enhanced by UniverSE++ showed decent restoration of the low-frequency harmonics, it lost many spectral details in the high-frequency regions, which significantly degraded the listening experience. These results suggest that baseline SE models are less effective in handling codec-induced compression distortions, likely due to the strong correlation between compressed and clean speech signals. In comparison, the proposed UDSE tackles this challenge from a discrete-domain perspective.
@@ IV-D3 Experimental Results for Mixed Distortions
In the three mixed distortion restoration tasks, our proposed UDSE had a significant advantage over all baseline models in both objective and subjective aspects ($p<0.01$), as shown in Tables I and II. For task DN+DR+BWE, UDSE outperformed UniverSE++ by more than 1-point in the UTMOS score. In addition, for task DN+DR+DC, UDSE achieved a 0.4-point higher OVRL score in DNSMOS and a 1.1-point higher UTMOS score compared to UniverSE++. Consistently, for task DN+PDR+CDR, UDSE again exceeded UniverSE++ by over 0.4-point in the UTMOS metric. Subjective evaluations showed that, in all three mixed-distortion tasks, the number of participants preferring UDSE-enhanced speech was at least double that of those favoring UniverSE++ — highlighting UDSE’s superior performance in managing complex distortions.
@@
The spectrograms shown in Figure 5 further provide a clear visual illustration of the effectiveness of the proposed UDSE in enhancing speech with mixed distortions. We can observe that for the DN+DR+BWE task (the first row in Figure 5), UDSE recovered the high-frequency components much more accurately, bringing the result closer to the ground-truth clean speech, while UniverSE++ struggled with recovering the high-frequency part. For the DN+DR+DC task (the second row in Figure 5), the speech enhanced by UniverSE++ lost a lot of detail in the high-frequency range and had a noticeable fundamental frequency error around 1 second, resulting in a poor listening experience. For the DN+PDR+CDR task (the last row in Figure 5), the spectrogram of the speech recovered by UDSE retained more details, while UniverSE++ produced significant high-frequency loss and poor spectral detail recovery, significantly impacting the listening experience. All the above experimental results demonstrate that the proposed UDSE can effectively mitigate various types of distortion in degraded speech, making it more suitable for real-world applications.